If you are standing 85 ft. from the base of a flag ploe, and determined that at ground level, you would need to look up at a 30 degree angle to see the top of the flag pole, how tall is the flag pole round the hight to tenths.
Answer by jojo14344(1513)

There you go, by trigonometric functions you can solve the height of the pole.
The distance of 85 ft from where you are standing to the height of the pole with an inclination of 30 degree is the hypotenuse.
Let x= height of flag pole.
by trigonometry, sin30 = x/hypotenuse
x=(sin30)(85)
x= 42.50 ft, height of flag pole. This is what you're looking for.
But you can also solve the distance from where you are standing to the pole.
By Pyth. theorem, let distance = d
(h)^2 = x^2 + d^2
85^2 = (42.50)^2 + d^2
d^2 = 7225-1806.25
d= sqrt 5418.75
d= 73.61 ft
Thank you,
Jojo